OpenStudy (1645323):

zero of this equation : Y=-6x^2+12x+21

OpenStudy (anonymous):

-6x^2+12x+21=0 solve the quadratic

OpenStudy (1645323):

-6x^2+12x=-21 what do i do from here?

OpenStudy (mathmale):

Which methods of finding the zeros of a function are you familiar with? Note: You may find it easier to find the zeros of Y=-6x^2+12x+21 if you notice that 3 is a factor common to all three terms and you factor that 3 out. What's left? Set this result = to 0 and find the zeros.

OpenStudy (anonymous):

take -3 common -3*(2x^2-4x-7)

OpenStudy (1645323):

thanks i got it now

OpenStudy (anonymous):

now either use perfect square method or factor method

OpenStudy (1645323):

the closest i get to zero is 1 and i use the number 3

OpenStudy (1645323):

@Zarkon

OpenStudy (anonymous):

2x^2-4x-7=0 \[\frac{ 4\pm \sqrt{4^2-4*(-7)*2} }{ 2*2 }\]

OpenStudy (1645323):

3.121320344

OpenStudy (anonymous):

\[\frac{ 4\pm \sqrt{16+56} }{ 4 }=\frac{ 4\pm \sqrt{72} }{ 4}\]

OpenStudy (anonymous):

sqrt72=8.48

OpenStudy (anonymous):

\[\frac{ 4+8.48 }{ 4 }=12.48/4=3.12\]

OpenStudy (1645323):

ok

OpenStudy (anonymous):

\[\frac{ 4-8.48 }{ 4 }=-4.48/4=-1.12\]

OpenStudy (anonymous):

@1645323